OpenStudy (anonymous):

Which is the 23rd term of the arithmetic series 5, 16, 27, ...? (2

OpenStudy (anonymous):

first term , a1 = 5 the difference between terms, d = 11 applying it into the 23rd term a23 = a1 + (n-1)d = 5 + 22*11 = 247

OpenStudy (anonymous):